find the equation of the line that passes through (-5,3) with its intercepts are numerically equal but opposite in direction.
Answer by nerdybill(7384)

find the equation of the line that passes through (-5,3) with its intercepts are numerically equal but opposite in direction.
.
Let x = x-intercept at (x,0)
then
-x = y-intercept at (0,-x)
.
slope of the two points (x,0) and (0,-x) is:
m = (y2-y1)/(x2-x1)
m = (-x-0)/(0-x)
m = (-x)/(-x)
m = 1
.
Use the slope (1) and the given point (-5,3) and plug into "point-slope" form:
y - y1 = m(x - x1)
y - 3 = 1(x - (-5))
y - 3 = 1(x + 5)
y - 3 = x + 5
y = x + 8 (this is what they're looking for)